Sol Summary – November 23rd

Sol Summary Report

November 23, 2017

Julielynn Wong, Crew 183 Mission Commander

Mission Status: Crew 183 received MDRS orientation training from MDRS Director, Shannon Rupert and EVA training from Crew 182 Geologist, Brandon Ferguson.

Sol Activity Summary: Yesterday, Crew 183 unloaded our equipment and personal items from our Tesla Model S vehicle.  Our Crew Astronomer and Engineer conducted night-time astral photography.  Our Greenhab Officer and Crew engineer assembled our portable off-grid power systems.

Today Crew 183 hiked around MDRS to observe the local geography.  We set up our Medical Maker Lab workstation on the first level of the habitat.  We have assembled all three of our humanitarian 3D printing systems (H3DPS-LM) which are operational.  We printed our dental tool file designed for the ISS and completed the print log.  Our Crew Engineer and Greenhab Officer have now built the systems to measure (i) power usage for each print and (ii) power generation by our bike and solar-powered 3D printing systems.  The Mission Commander prepared our breakfast and Thanksgiving dinner.

Look Ahead Plan: Preparing for sim on 0700 MST 2017-11-25 (Sol 3) and 0700 MST 2017-11-27 (Sol 5).

Anomalies in work:  Out of sim.

Weather:  Calm, sunny and warm.

Crew Physical Status:  Healthy

EVA: Crew Astronomer, Crew Engineer, and Greenhab Officer participated in an EVA led by our Crew Geologist.

Reports to be filed:

GreenHab Report
HSO Report
Operations Report
EVA Report
Astronomer Report
Support Requested: None
Julielynn Wong, MD, MPH
Crew 183 Mission Commander

EVA Report – November 23rd

EVA #1 Report

Today we conducted an out-of-sim EVA to the Dinosaur Quarry area, starting at approximately 3:00 PM. Ashton, Toluwa, Jordan, and I went on this EVA, each of us on our own ATV. The main purpose of today’s EVA was to drive the Blue 300 ATV longer than 5 km, so that its battery could recharge.

Before we set out, we did our ATV / Rover training, which lasted for about 20 minutes. Then we had to jumpstart the Blue 300 ATV, which had a low battery. By about 3:00 PM, we had the blue ATV running and ready to go. It took approximately 15-20 minutes for us to get there.

Once we arrived, I had a bit of a scare trying to restart the Blue ATV. It wouldn’t start at all! Then I realized that it was still in High. So I switched it to neutral and it started without any problem.

We spent over an hour in this area, following the canyon from the quarry site until we ended up in Lith Canyon, at the site where my EVA for crew 182 was conducted. We passed by one of the sites where I previously took regolith samples from. Just around the corner from this site was a large pile of petrified wood. Amazingly, these specimens look like fresh kindling wood, ready for the fireplace.

By around 4:30 PM, we made our way back to the ATVs, drove back to the hab, and returned by approximately 4:50 PM.

Have a great night Mission Support!

Prepared by Brandon Ferguson

EVA Officer

GreenHab Report – November 23rd

Sol 1 – 23/11/2017

Green Hab Officer Jordan Bowden

Average temperature:

Time 19:50
Temperature inside (°C) 19 C
Temperature outside (°C) 7 C
% Humidity 19 %

Changes to crops:  N/A – First day of mission

Daily water usage for crops: 3 Gallons

Time(s) of watering for crops:  19:50

Daily watering and amount of water used: 2.5

Narrative: We received our orientation late this morning, and as such I’ve only begun my work in the Greenhab. It’s an exciting place though: previous crew have left a growing crop of tomatoes, radishes, beans, and other crops. I was advised to use four gallons of water per day, but I only used 2.5 gallons, as the crops appeared to be well watered in the morning.

When I checked its status of the Greenhab at 19:50, the heater was not heating – not sure if it is supposed to be online at this time. It was working earlier today when checked at 12:00. In the end I sorted out that the thermostat was set to ~60 F. Please advise what the appropriate temperature should be.

I’m here mainly as a 3D Printing specialist – I’m wondering if there are any components to improve the work in the Greenhab that I could create?

Have a wonderful night everyone, I’m looking forward to another beautiful day on Mars.

All the best,

Jordan Bowden

Operations Report – November 23rd

Operations Report

Crew 183 Operations Report – 23 Nov 2017

SOL: 1

Name of person filing report:  Ashton Stoop

Non-nominal systems:

WATER HEATER: Water heater is still not installed. Hopefully this will be put into place soon.

Generator—  The generator has been turned on at 18:10.

READING FROM 23 NOV 2017

TIME

SOC READINGS

9:20

100%

12:20

100%

15:20

78%

18:10

62%

19:20

62%

20:05

62%

Diesel – 63 %

Oil Pressure: 47.5 PSI

Propane – 43% volume

Gasoline (5 Gallon containers for ATV) – 1 gallons.

Water (trailer) –  0 gallons

Water (static) – 40 gallons

Trailer to Static Pump used –   no

Water (loft) – Static to Loft Pump used –  yes

Water Meter:  47 gallons

Toilet tank emptied:  no

ATV’s Used: Yes

Oil Checked:  No

Atv Fuel Used Gals:  3  gallons

Tires Status: Perfect

# Hours the ATVs were Used today: 2 hours

Deimos used:  Yes

Hours: 0.5 hours

Beginning charge:  100%

Ending charge:  –

Currently charging: No

HabCar used? No

Notes and Comments:
The door to the Science Dome was not closing properly. The metal latch had been bent and was no longer sliding and latching the door. I was able to bend the metal back into place and the latch is now operational again.

We are having some issues with the power system. At 18:10 the SOC was at 62%. The generator was then turned on, but after 2.6 hours of using the generator, the SOC is still at 62%. My first thought is simply that it’s not reading correctly, but I am slightly concerned that it’s not charging properly.

Summary of GreenHab operations:

The GreenHab is looking in pretty good shape. There is a red wheelbarrow which has yet to be assembled. The axel is long enough to house the wheels, and all the parts are present, however there are protruding metal sleeves covering part of the axel. If this metal could be removed, then the wheels would easily slide on.

Summary of ScienceDome operations:

N/A

Summary of health and safety issues:

The carbon monoxide detector is not working, even with new batteries it won’t turn on.

Questions, concerns and requests to Mission Support:

-We are going to need gasoline for the ATVs ASAP.

-We need to figure out this whole power system.